A gas X has a vapour density of 32. It reacts with sodium hydroxide solution to form salt and water only. It decolourise acidified potassium permanganate solution and reacts with H2S to deposit sulphur. Identify the gas
1
Expert's answer
2020-05-03T15:11:50-0400

"vapour density = \frac{M_{gas}}{M_{H_2}}"

"M_{gas}=M_{H_2}\times (vapour density)=32\times 2.016=64.51 g/mol"


The formula of a gas is SO2


Reactions

1) "SO_2+2NaOH \rightarrow Na_2SO_3 + H_2O"

2) "5SO_2+2KMnO_4+2H_2O \rightarrow 2H_2SO_3 + 2MnSO_4+K_2SO_4"

3) "SO_2+2H_2S \rightarrow 3S+2H_2O"